OpenStudy (carlos41):

Hi, my name is Carlos. I do need help with the following: given the equation f(x)= 5x^2-4x, find the slope at (a, f(a)) and the equation at (2, f(2)) - The answer is: m = 10a – 4 and the equation at (2, f(2)) is : y=16x – 20. But I cannot get the same result using the the following: m= f(x) – f(a) / x – a So, using that formula, it begins like this: (5x^2-4x) – (5a^2-4a) / x – a .. . Any idea ? Thank you very much

OpenStudy (dumbcow):

to get the slope at point (a,f(a)) you need to take limit of "slope formula" as x approaches a: \[\lim_{x \rightarrow a} \frac{f(x) - f(a)}{x-a}\] this is easier to do if you make the limit approaching zero.... sub in a new variable h = x-a ...... a = x-h \[\rightarrow \lim_{h \rightarrow 0} \frac{f(x) - f(x-h)}{h}\] \[\lim_{h \rightarrow 0} \frac{(5x^2 -4x) - (5(x-h)^2 -4(x-h))}{h}\] \[\lim_{h \rightarrow 0} \frac{10xh - 5h^2 -4h}{h}\] Factor out an "h" \[\lim_{h \rightarrow 0} (10x -4 - 5h) = 10x - 4\] we are evaluating slope at x = a, \[m = 10a - 4\]

OpenStudy (dumbcow):

equation of line comes from using point-slope form \[y - f(2) = m (x - 2)\] \[y = mx -2m + f(2)\]

OpenStudy (carlos41):

dumbcow.. . THANK YOU VERY MUCH !! Appreciated